The heights of 500 students are normally distributed with a mean of 144.5 centimeters and a variance of 36 centimeters. Assuming that the heights are recorded to the nearest half-centimeter, find the Probability.
(a) Less than 130.0 centimeters?
(b) Between 151.5 and 182.0 centimeters inclusive?
(c) Greater than or equal to 150.0 centimeters?
(d) P(x > x0) = 0.6
1
Expert's answer
2020-06-30T18:42:16-0400

Since heights are normally distributed, the probability "P(x\leq x_0)" is given by (see https://en.wikipedia.org/wiki/Normal_distribution#Standard_normal_distribution):

"P(x\leq x_0)=\frac{1}{\sigma\sqrt{2\pi}}\int_{-\infty}^{x_0}e^{-\frac12(\frac{x-\mu}{\sigma})^2}dx"

with "\mu=144.5" , "\sigma=6" . In a similar way, we can express the probabilities "P(x_1\leq x\leq x_2)"

and "P(x\geq x_3)" for real "x_0,x_1,x_2,x_3" by changing the boundaries in the integral. We obtained the following answers:

(i) "P(x\leq 130)=\frac{1}{6\sqrt{2\pi}}\int_{-\infty}^{130}e^{-\frac12(\frac{x-144.5}{6})^2}dx\approx0.0078"

(ii) "P(151.5\leq x\leq 182)=\frac{1}{6\sqrt{2\pi}}\int_{151.5}^{182}e^{-\frac12(\frac{x-144.5}{6})^2}dx\approx0.1217"

(iii)"P(x\geq 150)=\frac{1}{6\sqrt{2\pi}}\int_{150}^{+\infty}e^{-\frac12(\frac{x-144.5}{6})^2}dx\approx0.1797"

(iv) The aim is to find "F^{-1}(0.6)" , where "F(y)=P(x> y)=\frac{1}{6\sqrt{2\pi}}\int_y^{+\infty}e^{-\frac12(\frac{x-144.5}{6})^2}dx"

For "y\approx143.0" (it was rounded to 1 decimal place) we get that "F(143.0)\approx0.5987"

For the computations we used the following program written in Anaconda (https://www.anaconda.com/products/individual):


from scipy import integrate

import numpy as np

import math


func = lambda x:(1/(6*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-144.5)/6)*((x-144.5)/6))

Pr1 = integrate.quad(func, -np.Infinity, 130)

Pr2 = integrate.quad(func, 151.5, 182)

Pr3 = integrate.quad(func, 150, +np.Infinity)


print(Pr1)

print(Pr2)

print(Pr3)


For the last task we used the fact that "F(y)" decreases and substituted different values of "y" until we got the values near 0.6.

Answers: (i) 0.0078 (ii) 0.1217 (iii) 0.1797 (iv) 143.0; the answers (i)-(iii) are rounded to 4 decimal places; (iv) is rounded to 1 decimal place.